Find all the cube roots of


1
Expert's answer
2021-08-24T16:00:27-0400

Find all the cube roots of "1+i."

The polar form of "1+i" is "\sqrt{2}(\cos\dfrac{\pi}{4}+i\sin \dfrac{\pi}{4})."

According to the De Moivre's Formula, all "n"-th roots of a complex number "r(\cos(\theta)+i\sin(\theta))" are given by "r^{1/n}(\cos\dfrac{\theta+2\pi k}{n}+i\sin \dfrac{\theta+2\pi k}{n}), k=0, 1,...,n-1."

We have that "r=\sqrt{2}, \theta=\dfrac{\pi}{4}, n=3."

"k=0"


"2^{1/6}(\cos\dfrac{\pi }{12}+i\sin\dfrac{\pi }{12})"


"k=1"


"2^{1/6}(\cos(\dfrac{\pi }{12}+\dfrac{2\pi}{3})+i\sin(\dfrac{\pi }{12}+\dfrac{2\pi}{3}))"

"=2^{1/6}(\cos(\dfrac{3\pi }{4})+i\sin(\dfrac{3\pi }{4}))"

"k=2"

"2^{1/6}(\cos(\dfrac{\pi }{12}+\dfrac{4\pi}{3})+i\sin(\dfrac{\pi }{12}+\dfrac{4\pi}{3}))"

"=2^{1/6}(\cos(\dfrac{17\pi }{12})+i\sin(\dfrac{17\pi }{12}))""=-2^{1/6}(\cos(\dfrac{5\pi }{12})+i\sin(\dfrac{5\pi }{12}))"